# Adaptive Cache Enhancement for Test-Time Adaptation of Vision-Language Models

### v1

@@ Introduction
Following prior research (Karmanov et al. 2024a; Zhang et al. 2024a), we evaluate our method in two main benchmarking scenarios: cross-dataset generalization and robustness to natural distribution shifts. (1) For the cross-dataset generalization scenario, we carry out evaluations across 10 varied recognition datasets, including FGVC Aircraft (Maji et al. 2013a), Caltech101 (Fei-Fei, Fergus, and Perona 2004), Stanford Cars (Krause et al. 2013a), DTD (Cimpoi et al. 2013), EuroSAT (Helber et al. 2019), Flowers102 (Nilsback and Zisserman 2008a), Food101 (Bossard, Guillaumin, and Van Gool 2014), Oxford Pets (Parkhi et al. 2012a), SUN397 (Xiao et al. 2010a), and UCF101 (Soomro 2012). These datasets collectively offer a comprehensive benchmark to assess the adaptability and generality of models across different datasets. (2) To evaluate the robustness of our method against natural distribution shifts, we test our method with ImageNet (Deng et al. 2009) and its out-of-distribution variants, including ImageNet-A (Hendrycks et al. 2021b) and ImageNet-V2 (Recht et al. 2019). This part of the evaluation assesses our method’s robustness under varying distribution shifts. For datasets with particularly large test sets, e.g., Food101, ImageNet, SUN397, ImageNet-V2, and ImageNet-A, we sample 2,000 test instances for evaluation, adhering to the experimental protocol in DiffTPT (Feng et al. 2023c).
@@
Following previous studies (Karmanov et al. 2024a; Zhang et al. 2024a), we select ResNet-50 and ViT-B/16 as visual encoders for CLIP, employing hand-crafted prompts as described in the supplementary material. In line with the method by (Shu et al. 2022a), we create 63 augmented views per test image. The AdamW optimizer is used with the learning rate set to 0.0005 for the residual parameters in one iteration. Hyperparameters are set as follows: $\alpha=6$, $\beta=5.0$, $\delta=0.95$, $\gamma=0.02$, $\lambda=0.5$, and queue size $M=16$ (since ACE shows resilience to noisy pseudo-labels and benefits from additional samples for class distribution modeling). All experiments are conducted on a single NVIDIA A100 GPU with 80GB of memory.
Results
@@
Our method reliably excels in adapting to varied image datasets with natural shifts, as highlighted in Table 1. Compared to TDA (Karmanov et al. 2024a) and DPE (Zhang et al. 2024d), both our probability and entropy variants outperform them across 4 out of 5 datasets with notable differences, tested across ResNet-50 and ViT-B/16 backbones. Specifically, our ACE achieves top OOD average accuracies of 47.79% with ResNet-50 and 64.63% with ViT-B/16. While DiffTPT (Feng et al. 2023a) occasionally surpasses our method, we found this is due to the limited use of data augmentation in our experiments. Our method also outperforms the few-shot learning technique CoOp (Zhou et al. 2021), underscoring our method’s robust ability to adapt VLMs to OOD scenarios.
@@
In Table 2, we evaluate the generality of our ACE-Probability and ACE-Entropy variants across 10 diverse fine-grained datasets, and compare them with state-of-the-art approaches. Due to significant distributional differences among these datasets, performance varies considerably. However, our methods achieve average accuracy improvements of 1.65% and 0.82% over the best-performing methods on the CLIP-ResNet-50 and ViT-B/16 backbones, respectively, surpassing competitors in 6 out of 10 datasets. Notably, on the challenging DTD dataset, characterized by complex texture categories, our approach consistently outperforms prior works across both CLIP backbones by a large margin. Additionally, our approach significantly outperforms the few-shot prompt learning method CoOp (Zhou et al. 2021) in all datasets. These findings highlight the generality and adaptability of our approach in test-time domain transfer, making it highly suitable for real-world applications.
Ablation study
Cache size and zero-shot statistics.
@@
Strat. Cache size Caltech DTD Flower Pets UCF
@@
Visualization of cache features using t-SNE
@@
Figure 3 displays t-SNE (van der Maaten and Hinton 2008) of the cached image features (with cache size $M=6$) for both DPE (Zhang et al. 2024d) and our method on the Food101 (Bossard, Guillaumin, and Van Gool 2014) dataset. Features from 10% of randomly selected classes are shown in different colors, while all others are indicated in gray. The visualizations demonstrate that our priority queue approach successfully gathers high-confidence samples, leading to increasingly representative visual prototypes over time.
@@
Figure 3: t-SNE (van der Maaten and Hinton 2008) visualizations of the stored image features in the cache between DPE (Zhang et al. 2024d) and ACE-Entropy.
Cache reliability
@@
Figure 4 visualizes the cache accuracy and overall classification performance of TDA, DPE, and our ACE across four fine-grained datasets. ACE demonstrates exceptional results, achieving an accuracy of 68.55% on Stanford Cars, surpassing TDA by 1.27% and DPE by 1.24%, and 77.30% on Oxford Flowers, outperforming TDA by 5.88% and DPE by 2.07%. These gains are complemented by ACE’s consistent superiority in cache accuracy across the datasets, reinforcing its ability in enhancing classification power while sustaining reliable and flexible decision boundaries via class-wise thresholding adaptation.
@@
Figure 4: Comparison of cache accuracy vs test accuracy over four datasets between TDA (Karmanov et al. 2024a), DPE (Zhang et al. 2024d), and our proposed ACE-Entropy.
Conclusion
@@ Appendix A Appendix
Appendix B Detailed Dataset Information
@@
Caltech101 (Fei-Fei, Fergus, and Perona 2007) 100 4,128 1,649 2,465 Object recognition
DTD (Cimpoi et al. 2014) 47 2,820 1,128 1,692 Texture recognition
EuroSAT (Helber et al. 2019) 10 13,500 5,400 8,100 Satellite image recognition
FGVCAircraft (Maji et al. 2013b) 100 3,334 3,333 3,333 Fine-grained aircraft recognition
Flowers102 (Nilsback and Zisserman 2008b) 102 4,093 1,633 2,463 Fine-grained flowers recognition
Food101 (Bossard, Guillaumin, and Van Gool 2014) 101 50,500 20,200 30,300 Fine-grained food recognition
ImageNet (Deng et al. 2009) 1,000 1.28M - 50,000 Object recognition
OxfordPets (Parkhi et al. 2012b) 37 2,944 736 3,669 Fine-grained pets recognition
StanfordCars (Krause et al. 2013b) 196 6,509 1,635 8,041 Fine-grained car recognition
SUN397 (Xiao et al. 2010b) 397 15,880 3,970 19,850 Scene recognition
UCF101 (Soomro, Zamir, and Shah 2012) 101 7,639 1,898 3,783 Action recognition
ImageNet-V2 (Recht et al. 2019) 1,000 - - 10,000 Robustness of collocation
ImageNet-Sketch (Wang et al. 2019) 1,000 - - 50,889 Robustness of sketch domain
ImageNet-A (Hendrycks et al. 2021b) 200 - - 7,500 Robustness of adversarial attack
ImageNet-R (Hendrycks et al. 2021a) 200 - - 30,000 Robustness of multi-domains
@@
Appendix C Textual Prompts Used in Experiments
@@
ImageNet (Deng et al. 2009) “a bad photo of the {CLASS}.”
ImageNet-V2 (Recht et al. 2019) “a origami {CLASS}.”
ImageNet-Sketch (Wang et al. 2019) “a photo of the large {CLASS}.”
ImageNet-A (Hendrycks et al. 2021b) “a {CLASS} in a video game.”
ImageNet-R (Hendrycks et al. 2021a) “art of the {CLASS}.”
@@
Caltech101 (Fei-Fei, Fergus, and Perona 2007) “a photo of a {CLASS}.”
DTD (Cimpoi et al. 2014) “{CLASS} texture.”
EuroSAT (Helber et al. 2019) “a centered satellite photo of {CLASS}.”
FGVCAircraft (Maji et al. 2013b) “a photo of a {CLASS}, a type of aircraft.”
Flowers102 (Nilsback and Zisserman 2008b) “a photo of a {CLASS}, a type of flower.”
Food101 (Bossard, Guillaumin, and Van Gool 2014) “a photo of {CLASS}, a type of food.”
OxfordPets (Parkhi et al. 2012b) “a photo of a {CLASS}, a type of pet.”
StanfordCars (Krause et al. 2013b) “a photo of a {CLASS}.”
SUN397 (Xiao et al. 2010b) “a photo of a {CLASS}.”
UCF101 (Soomro, Zamir, and Shah 2012) “a photo of a person doing {CLASS}.”
Table 5: Textual prompts used in experiments. In addition to these prompts, we also employ CuPL (Pratt et al. 2023) prompts to further enhance performance.
Appendix D Full results of ablation study

### v2

@@ 4.1 Experimental Setup
Following prior research [tda, dpe], we evaluate our method in two main benchmarking scenarios: cross-dataset generalization and robustness to natural distribution shifts. (1) For the cross-dataset generalization scenario, we carry out evaluations across 10 varied recognition datasets, including FGVC Aircraft [aircraft], Caltech101 [caltech101], Stanford Cars [cars], DTD [Dtd], EuroSAT [helber2019eurosat], Flowers102 [oxfordflower], Food101 [bossard2014food], Oxford Pets [oxford_pets], SUN397 [sun397], and UCF101 [ucf101]. These datasets collectively offer a comprehensive benchmark to assess the adaptability and generality of models across different datasets. (2) To evaluate the robustness of our method against natural distribution shifts, we test our method with ImageNet [deng2009imagenet] and its out-of-distribution variants, including ImageNet-A [hendrycks2021natural] and ImageNet-V2 [recht2019imagenet]. This part of the evaluation assesses our method’s robustness under varying distribution shifts. For datasets with particularly large test sets, e.g., Food101, ImageNet, SUN397, ImageNet-V2, and ImageNet-A, we sample 2,000 test instances for evaluation, adhering to the experimental protocol in DiffTPT [Feng2023DiverseDA].
@@
Following previous studies [tda, dpe], we select ResNet-50 and ViT-B/16 as visual encoders for CLIP, employing hand-crafted prompts as described in the supplementary material. In line with the method by [tpt], we create 63 augmented views per test image. The AdamW optimizer is used with the learning rate set to 0.0005 for the residual parameters in one iteration. Hyperparameters are set as follows: $\alpha=6$, $\beta=5.0$, $\delta=0.95$, $\gamma=0.02$, $\lambda=0.5$, and queue size $M=16$ (since ACE shows resilience to noisy pseudo-labels and benefits from additional samples for class distribution modeling). All experiments are conducted on a single NVIDIA A100 GPU with 40GB of memory.
4.2 Results
@@
Our method reliably excels in adapting to varied image datasets with natural shifts, as highlighted in Table 2. Compared to TDA [tda] and DPE [zhang2024dual], both our probability and entropy variants outperform them across 4 out of 5 datasets with notable differences, tested across ResNet-50 and ViT-B/16 backbones. Specifically, our ACE achieves top OOD average accuracies of 49.38% with ResNet-50 and 66.05% with ViT-B/16. While CRG [zhai2025mitigatingcachenoisetesttime] occasionally surpasses our method, we found this is due to the limited use of data augmentation in our experiments. Our method also outperforms the few-shot learning technique CoOp [CoOp], underscoring our method’s robust ability to adapt VLMs to OOD scenarios.
@@
In Table 3, we evaluate the generality of our ACE-Probability and ACE-Entropy variants across 10 diverse fine-grained datasets, and compare them with state-of-the-art approaches. Due to significant distributional differences among these datasets, performance varies considerably. However, our methods achieve average accuracy improvements of 1.65% and 0.82% over the best-performing methods on the CLIP-ResNet-50 and ViT-B/16 backbones, respectively, surpassing competitors in 6 out of 10 datasets. Notably, on the challenging DTD dataset, characterized by complex texture categories, our approach consistently outperforms prior works across both CLIP backbones by a large margin. Additionally, our approach significantly outperforms the few-shot prompt learning method CoOp [CoOp] in all datasets. These findings highlight the generality and adaptability of our approach in test-time domain transfer, making it highly suitable for real-world applications.
4.3 Ablation study
4.3.1 Cache size and zero-shot statistics.
@@ 4.3 Ablation study
Strategy Cache size Caltech DTD Flower Pets UCF
@@
4.3.2 Visualization of cache features using t-SNE
@@
Figure 3 displays t-SNE [van2008visualizing] of the cached image features (with cache size $M=6$) for both DPE [zhang2024dual] and our method on the Food101 [bossard2014food] dataset. Features from 10% of randomly selected classes are shown in different colors, while all others are indicated in gray. The visualizations demonstrate that our priority queue approach successfully gathers high-confidence samples, leading to increasingly representative visual prototypes over time.
@@
Figure 3: t-SNE [van2008visualizing] visualizations of the stored image features in the cache between DPE [zhang2024dual] and ACE-Entropy.
4.3.3 Efficiency Analysis
The efficiency analysis in the Table 5 shows the trade-off of different methods. Baseline zero-shot CLIP achieves 59.81% accuracy in just 9 minutes, serving as a reference for minimal overhead, while prompt-based approaches like TPT and DiffTPT deliver modest gains of +0.93% and +0.99% at significantly higher costs of 9 hours 15 minutes and 20 hours, respectively, due to backpropagation-intensive prompt optimization. Cache-based methods, including TDA (+1.54% in 1 hour 5 minutes), TPS (+1.66% in 55 minutes), and DPE (+3.60% in 1 hour 50 minutes), offer better efficiency through lightweight mechanisms. Notably, the proposed ACE variants stand out, with ACE-Probability attaining the highest accuracy of 64.41% (+4.60% gain) and ACE-Entropy closely following at 64.30% (+4.49% gain), both in 2 hours 40 minutes; this moderate runtime increase reflects ACE’s dynamic threshold refinement via exponential moving average and exploration updates, enabling superior robustness and generalization without prohibitive computational demands, positioning it as a state-of-the-art solution for practical out-of-distribution adaptation.
Method Testing Time Accuracy Gain
CLIP [radford2021learning] 9 min 59.81 -
TPT [shu2022test] 9 h 15 min 60.74 +0.93
DiffTPT [feng2023diverse] 20 h 00 min 60.80 +0.99
TDA [karmanov2024efficient] 1 h 5 min 61.35 +1.54
TPS [sui2024just] 55 min 61.47 +1.66
DPE 1 h 50 min 63.41 +3.60
ACE-Probability 2 h 40 min 64.41 4.60
ACE-Entropy 2 h 40 min 64.30 4.49
Table 5: Efficiency analysis of ACE vs other methods.
4.3.4 Cache reliability
@@
Figure 4 visualizes the cache accuracy and overall classification performance of TDA, DPE, and our ACE across four fine-grained datasets. ACE demonstrates exceptional results, achieving an accuracy of 68.55% on Stanford Cars, surpassing TDA by 1.27% and DPE by 1.24%, and 77.30% on Oxford Flowers, outperforming TDA by 5.88% and DPE by 2.07%. These gains are complemented by ACE’s consistent superiority in cache accuracy across the datasets, reinforcing its ability in enhancing classification power while sustaining reliable and flexible decision boundaries via class-wise thresholding adaptation. The synergy between these metrics underscores their complementary nature, enhancing overall performance.
@@
Figure 4: Comparison of cache accuracy vs. test accuracy over four datasets between TDA [tda], DPE [zhang2024dual], and our ACE-Entropy.
4.3.5 Efficiency Comparison
The inference efficiency and adaptation performance of mainstream test-time adaptation (TTA) approaches are comprehensively assessed on ImageNet [deng2009imagenet], with the outcomes summarized in Table 5. Compared to the zero-shot CLIP [radford2021learning] baseline, which achieves 59.81% top-1 accuracy within 11 minutes, ACE attains a markedly higher accuracy of 64.41% after 2 hours and 40 minutes of adaptation—representing an improvement of +4.60%. Although TPT [shu2022test] and DiffTPT [feng2023diverse] obtain similar accuracies of 60.74% and 60.80%, respectively, both require over 10 and 20 hours of computation, resulting in substantial computational overhead that constrains their practical applicability in efficiency-critical scenarios. TDA [karmanov2024efficient], while more efficient with a runtime of 32 minutes, achieves a lower accuracy of 61.35%. CRG [zhai2025mitigatingcachenoisetesttime] although achieves higher accuracy of 65.26%, it significantly requires much more time to run, result in a 3 hours running. Overall, ACE delivers the highest accuracy among all evaluated methods while maintaining a moderate adaptation time. Its adaptive components—such as dynamic curriculum thresholding and online thresholding adaptation - appears to enable faster convergence and more robust performance. These findings indicate that ACE strikes a more effective balance between adaptation speed and accuracy compared to alternative techniques, underscoring its methodological strengths in test-time adaptation.
5 Conclusion
@@ 6 Appendix
7 Detailed Dataset Information
@@
Caltech101 [fei2004learning] 100 4,128 1,649 2,465 Object recognition
DTD [cimpoi2014describing] 47 2,820 1,128 1,692 Texture recognition
EuroSAT [helber2019eurosat] 10 13,500 5,400 8,100 Satellite image recognition
FGVCAircraft [maji2013fine] 100 3,334 3,333 3,333 Fine-grained aircraft recognition
Flowers102 [nilsback2008automated] 102 4,093 1,633 2,463 Fine-grained flowers recognition
Food101 [bossard2014food] 101 50,500 20,200 30,300 Fine-grained food recognition
ImageNet [deng2009imagenet] 1,000 1.28M - 50,000 Object recognition
OxfordPets [parkhi2012cats] 37 2,944 736 3,669 Fine-grained pets recognition
StanfordCars [krause20133d] 196 6,509 1,635 8,041 Fine-grained car recognition
SUN397 [xiao2010sun] 397 15,880 3,970 19,850 Scene recognition
UCF101 [soomro2012ucf101] 101 7,639 1,898 3,783 Action recognition
ImageNet-V2 [recht2019imagenet] 1,000 - - 10,000 Robustness of collocation
ImageNet-Sketch [wang2019learning] 1,000 - - 50,889 Robustness of sketch domain
ImageNet-A [hendrycks2021natural] 200 - - 7,500 Robustness of adversarial attack
ImageNet-R [hendrycks2021many] 200 - - 30,000 Robustness of multi-domains
@@
8 Textual Prompts Used in Experiments
@@
ImageNet [deng2009imagenet] “a bad photo of the {CLASS}.”
ImageNet-V2 [recht2019imagenet] “a origami {CLASS}.”
ImageNet-Sketch [wang2019learning] “a photo of the large {CLASS}.”
ImageNet-A [hendrycks2021natural] “a {CLASS} in a video game.”
ImageNet-R [hendrycks2021many] “art of the {CLASS}.”
@@
Caltech101 [fei2004learning] “a photo of a {CLASS}.”
DTD [cimpoi2014describing] “{CLASS} texture.”
EuroSAT [helber2019eurosat] “a centered satellite photo of {CLASS}.”
FGVCAircraft [maji2013fine] “a photo of a {CLASS}, a type of aircraft.”
Flowers102 [nilsback2008automated] “a photo of a {CLASS}, a type of flower.”
Food101 [bossard2014food] “a photo of {CLASS}, a type of food.”
OxfordPets [parkhi2012cats] “a photo of a {CLASS}, a type of pet.”
StanfordCars [krause20133d] “a photo of a {CLASS}.”
SUN397 [xiao2010sun] “a photo of a {CLASS}.”
UCF101 [soomro2012ucf101] “a photo of a person doing {CLASS}.”
Table 7: Textual prompts used in experiments. In addition to these prompts, we also employ CuPL [pratt2023does] prompts to further enhance performance.
9 Full results of ablation study
@@
10 Supplementary for different seeds
We report the complete results of ImageNet with 3 different seeds in Table 9. Since we introduce the Nnever and rarely seen adaptation, and ACE is only applied during testing, the sequence of testing samples could affect the performance. Here, as we can see in Table 9, the standard deviation between datasets is stable, showing that ACE performance is not affected by randomness.
CLIP-ResNet-50 [radford2021learning] 58.16 21.83 51.41 56.15 33.37 44.18 40.69
Ensemble 59.81 23.24 52.91 60.72 35.48 46.43 43.09
CoOp [zhou2022learning] 63.33 23.06 55.40 56.60 34.67 46.61 42.43
TPT [shu2022test] 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [feng2023diverse] 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [karmanov2024efficient] 61.35 30.29 55.54 62.58 38.12 49.58 46.63
TPS [sui2024just] 61.47 30.48 54.96 62.87 37.14 49.38 46.36
DMN-ZS [zhang2024dual] 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE [dpe] 63.41 30.15 56.72 63.72 40.03 50.81 47.66
CRG [zhai2025mitigatingcachenoisetesttime] 65.26 29.69 56.07 - - - -
[…]
CLIP-ViT-B/16 [radford2021learning] 66.73 47.87 60.86 73.98 46.09 59.11 57.20
Ensemble 68.34 49.89 61.88 77.65 48.24 61.20 59.42
CoOp [zhou2022learning] 71.51 49.71 64.20 75.21 47.99 61.72 59.28
TPT [shu2022test] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [feng2023diverse] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [karmanov2024efficient] 69.51 60.11 64.67 80.24 50.54 65.01 63.89
TPS [sui2024just] 70.19 60.08 64.73 80.27 49.95 65.04 63.76
DMN-ZS [zhang2024dual] 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE [dpe] 71.91 59.63 65.44 80.40 52.26 65.93 64.43
CRG [zhai2025mitigatingcachenoisetesttime] 75.01 63.67 64.66 - - - -
[…]
Caltech +0.20
DTD +1.10
Flowers +0.80
Pets +0.1
UCF101 +1.2
Average +0.68
[…]
Caltech +0.30
DTD +1.70
Flower +1.50
Pets +0.70
UCF +1.40
Average +1.12
[…]
Caltech +0.45
DTD +2.10
Flower +2.30
Pets +1.50
UCF +2.80
Average +1.83
[…]
Caltech +0.32
DTD +2.18
Flower +2.40
Pets +0.33
UCF +3.18
Average +1.68
Table 10: Ablation study of each of the thresholding components.
11 Ablation of Thresholding Components
To better understand the contribution of each part of our adaptive thresholding design, we decompose ACE into four sub-components corresponding to the individual tables in Fig. 10. Each sub-table (a)–(d) reports accuracy improvements (%) over the DPE baseline across the five datasets used in our main evaluation.
(a) Curriculum Thresholding (CT).
Table LABEL:tab:ct_only isolates the effect of CT alone. By allowing class-wise thresholds to evolve based on the model’s progress, CT already provides substantial improvements over DPE across all datasets, most notably on DTD and Flowers102. These results demonstrate that dynamically adjusting thresholds is fundamental to increasing cache purity and stabilizing early adaptation.
(b) EMA-Based Refinement.
Adding EMA smoothing on top of CT (Table LABEL:tab:ema_only) further improves accuracy on every dataset. EMA suppresses noisy fluctuations in threshold values and stabilizes pseudo-label evolution, leading to a consistent gain of roughly +1% over CT alone. This confirms that temporal smoothing is crucial for reliable threshold updates in an online setting.
(c) Exploration for Rare Classes.
Table LABEL:tab:exploration evaluates the effect of enabling our exploration mechanism. By decreasing thresholds for underrepresented classes, exploration increases cache diversity and improves performance on datasets with high intra-class variance or long-tail structure. The gains over (b) are the largest among all incremental components, highlighting exploration as a key driver for handling rare or difficult categories.
(d) Zero-Shot Initialization (ZS-Init).
Finally, Table LABEL:tab:zs_init shows the effect of adding zero-shot threshold initialization. ZS-Init provides a strong, distribution-aware starting point for each class, preventing early miscalibration and significantly improving early adaptation stability. When combined with CT, EMA, and exploration, ZS-Init produces the highest overall accuracy improvement, confirming its role in accelerating convergence and improving cache reliability.
Summary.
Across all four sub-components, each contributes positively and incrementally to the final performance of ACE. The progression from (a) to (d) demonstrates that curriculum scheduling, temporal smoothing, exploration, and zero-shot initialization are individually beneficial and complementary when combined.